A 10 Kg block kept on a double inclined plane is pulled by a string with a 200N force. A 10N force is also applied on the 10 Kg block(angle of inclination for 10 kg block-30 degree). Find: a)tension in string, b)net force on pulley exerted by string

Grade:11

1 Answers

Jatin
5 years ago
Here we have given that the tension on the string is 200N because there is no block on other side that force means  that 200N only will be the tension force. 
now we will make the equation on 10kg box  
T-10+Mgsin30°=10a
T-10+100 sin30°
T-60=10a........(1)
we are given that the tension is 200N
200-60=10a
140=10a 
a=14m/s